OpenStudy (anonymous):

Which correctly describes the roots of the following cubic equation? x^3-3x^2+4x-12=0 a) three real roots each with a different value b) one real root and two complexroots c) three real roots, two of which are equal in value d) two real roots and one complex root

OpenStudy (anonymous):

I know that the factors of 12 are 1, 2, 3, 4, 6, and 12

OpenStudy (anonymous):

Therefore, the possible roots are 1, 2, 3, 4, 6, and 12

OpenStudy (anonymous):

does it help to know that it is \[(x-3)(x^2+4)\]?

OpenStudy (anonymous):

actually there are more possible rational zeros, also the negative of what you wrote above a rather large total, and there is no guarantee that the zeros are even rational, although they are in this example

OpenStudy (anonymous):

(x-3)(x^2+4) Okay, so for (x-3), x=3?

OpenStudy (anonymous):

Because x^2+4=0 creates x^2=-4, does this mean that there is one real root and two complex roots?

OpenStudy (anonymous):

@jdoe0001